OpenStudy (anonymous):

For the function f(x) = sin(x), on the interval [0, pi], which of the following is true?

OpenStudy (anonymous):

For f(x) = sin(x), the Mean Value Theorem states that for an interval [0, pi], there is a point x = c in (0,)where f(c) is equal to the average of f(b) and f(a). For f(x) = sin(x), The Mean Value Theorem does not apply. For f(x) = sin(x), The Mean Value Theorem tells you how to find a value x = c in (a, b) such that the instantaneous rate of change at c is equal to the average rate of change between a and b. For f(x) = sin(x), the Mean Value Theorem states that for an interval [0,pi ], there is a point x = c in (0, ) where the instantaneous rate of change at x = c is equal to the average rate of change between a and b.